OpenStudy (anonymous):

what is the square root of 100

OpenStudy (solomonzelman):

Okay, you know that \(\large\color{black}{ 10 \times 10=100 }\) ?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

can you tell me the answer based on that, or need a couple of examples?

OpenStudy (solomonzelman):

I will show you a couple of examples. ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ \(\large\color{black}{ \sqrt{81}=?}\) since, \(\large\color{black}{ 9 \times 9=81 }\) therefore: \(\large\color{black}{ \sqrt{81}=9}\) ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ \(\large\color{black}{ \sqrt{64}=?}\) since, \(\large\color{black}{ 8 \times8=64 }\) therefore: \(\large\color{black}{ \sqrt{64}=8}\) ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ \(\large\color{black}{ \sqrt{49}=?}\) since, \(\large\color{black}{ 7 \times 7=49 }\) therefore: \(\large\color{black}{ \sqrt{49}=7}\)

OpenStudy (solomonzelman):

~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ \(\large\color{black}{ \sqrt{100}=?}\) since, \(\large\color{black}{ 10 \times 10=100 }\) therefore: \(\large\color{black}{ \text{_____}}\)

OpenStudy (solomonzelman):

can you finish the last statement for me?

OpenStudy (anonymous):

you right it 's 10

OpenStudy (solomonzelman):

yes